OpenStudy (anonymous):

Solve the equation. 15/m - m + 8 = 10

OpenStudy (anonymous):

we can say : 15/m-m=10-8 k ?:)

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

you have to simplify the equation, and then factorise it at the end to get your answers

OpenStudy (anonymous):

15/m-m=10-8 15/m-m^2/m=2 15-m^2/m=2 Factor from m : m(m-15)/m=2 have : m-15=2 m=15+2 m=17 Got it ?:)

OpenStudy (anonymous):

lol i dont think thats it

OpenStudy (anonymous):

Why ? What is your way ?:)

OpenStudy (anonymous):

the answer is: (m+5)(m-3) which gives m = -5 or m = 3

OpenStudy (anonymous):

you can check these types of questions by subbing in your answers back into the original question, and is the left hand side and right hand side of the = sign are the same then you are right

OpenStudy (anonymous):

I want your way !

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

let me know if u need the working out part :)

OpenStudy (anonymous):

I do, just so I know I understand it cuz I have to do another problem like it

OpenStudy (anonymous):

15/m-m=10-8 (15/m) - m = 2 (15/M) - M - 2 = 0 now we are going to multiply all 3 terms by 'm' (15M/M) - m^2 - 2M = 0 so i will just re arrange it by taking everything to the otherside to make em positive numbers m^2 +2M - 15 =0 Have you seen equations like this be4? i

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so u know how to factorise them?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

good :P

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

anytime